OpenStudy (anonymous):

let f(x)=-4x+7 and g(x)=10x-6 find (f(g)) A. -40x+24 B. -40x+31 C. -40x+64 D. -40x+70 @joftheworld I would of replied but it wouldnt let me send you a message because you only accept messages from people you fanned lol sorry :(

OpenStudy (anonymous):

Know what (f(g)) means?

OpenStudy (anonymous):

composition of functions right?

OpenStudy (anonymous):

Yes.

OpenStudy (joftheworld):

yes composite functions

OpenStudy (anonymous):

So you are putting the function for g(x) inside of f(x).

OpenStudy (joftheworld):

in order to do that you have to find f(g)

OpenStudy (joftheworld):

is there anymore info on this problem ?

OpenStudy (anonymous):

No there is not anything else thats all it says

OpenStudy (anonymous):

You do not need more. You are given f(x) and g(x). You just need to replace the x in f(x) with the function part of g(x) and solve.

OpenStudy (anonymous):

When you put g(x) into f(x) what do you get?

OpenStudy (anonymous):

Here are some examples: m(x) = 2x+5 n(x) = 3x-1 m(n(x)) = 2(3x-1)+5 m(n(x)) = 6x-2+5 m(n(x)) = 6x+3 n(m(x)) = 3(2x+5)-1 n(m(x)) = 6x+15-1 n(m(x)) = 6x+14 You can see that all I am doing is putting one function into the other distributing and simplifying.